Six hundred liters of a solution was available, but the solution was 70% alcohol. Dan needed a solution of 40% alcohol. How many liters of alcohol had to be extracted so that the solution would be 40% alcohol?
Answer by ankor@dixie-net.com(22740)

Six hundred liters of a solution was available, but the solution was 70% alcohol.
Dan needed a solution of 40% alcohol.
How many liters of alcohol had to be extracted so that the solution would be 40% alcohol?
:
Let x = amt of alcohol to removed
.7(600) - x = .4(600-x)
420 - x = 240 - .4x
420 - 240 = -.4x + x
180 = .6x
x = 180/.6
x = 300 liters of alcohol